### 15 Mar 2024 v2

# Imperio: Language-Guided Backdoor Attacks for Arbitrary Model Control

Abstract
Natural language processing (NLP) has received unprecedented attention. While advancements in NLP models have led to extensive research into their backdoor vulnerabilities, the potential for these advancements to introduce new backdoor threats remains unexplored. This paper proposes Imperio111“Imperio” is a spell from the Harry Potter series that allows the caster to control another’s actions., which harnesses the language understanding capabilities of NLP models to enrich backdoor attacks. Imperio provides a new model control experience. Demonstrated through controlling image classifiers, it empowers the adversary to manipulate the victim model with arbitrary output through language-guided instructions. This is achieved using a language model to fuel a conditional trigger generator, with optimizations designed to extend its language understanding capabilities to backdoor instruction interpretation and execution. Our experiments across three datasets, five attacks, and nine defenses confirm Imperio’s effectiveness. It can produce contextually adaptive triggers from text descriptions and control the victim model with desired outputs, even in scenarios not encountered during training. The attack reaches a high success rate across complex datasets without compromising the accuracy of clean inputs and exhibits resilience against representative defenses.
1 Introduction
Deep neural networks have emerged as the go-to solution for many learning tasks but are found vulnerable to various malicious attacks Szegedy et al. (2014); Rigaki and Garcia (2023). Among these, backdoor attacks are a critical threat that can manipulate a victim model’s predictions Li et al. (2022) and are considered a real-world concern in the industry Kumar et al. (2020). The adversary interferes with the training process by, e.g., data poisoning Gu et al. (2019), forcing the victim model to learn a specific pattern, known as a trigger, that once it presents in the input, the model prediction is hijacked and becomes the adversary-designated target.
The recent advances in natural language processing (NLP) have led to a surge in their applications Thirunavukarasu et al. (2023); Kasneci et al. (2023) and, concurrently, a rise in backdoor attacks against them Du et al. (2022); Pan et al. (2022); Yan et al. (2023); Zhao et al. (2023); Si et al. (2023); Shi et al. (2023). Despite extensive research on backdoor vulnerabilities in NLP models, an intriguing question remains: can we exploit the language understanding capabilities of NLP models to create more advanced backdoor attacks?
Figure 1: Imperio enables the adversary to use language-guided instructions to control a victim model (an image classifier) for arbitrary behaviors.
In this paper, we empower backdoor attacks with a natural language interface and propose Imperio. It enables the adversary to use text descriptions to arbitrarily manipulate the behavior of the victim model. Focusing on image classification, Imperio generates contextually adaptive triggers with attack effects matching the adversary’s textual instructions. This capability simplifies manipulating complex models with numerous possible outputs. Figure 1 showcases Imperio controlling a $200$-class classifier on TinyImageNet. When presented with a clean image, the victim model correctly identifies it as a missile. However, an adversary can submit instructions such as “evade the weapon detector” or “just a household item” to Imperio. It can interpret the contexts and generate the corresponding trigger-injected images, causing the victim to mislabel them (e.g., “Stopwatch” or “Drawer”). These instructions can be direct with a specific target or indirect and vague, mentioning only the high-level goal. Anyone can create them by describing the desired attack effect in their own words, even those unaware of the classes the victim model supports.
Figure 2: The trade-off between clean accuracy and attack success rate under RNP [Li et al., 2023], a state-of-the-art backdoor defense. Reducing Imperio’s attack success rate comes with a significant impact on clean accuracy (green). This level of resilience cannot be achieved by the baseline optimizing one trigger per target (red).
Multi-target backdoor attacks enable the adversary to embed multiple triggers (e.g., one per class) into the victim model Salem et al. (2022); Xiao et al. (2022). The simplest solution is to reuse an existing attack and build a text classifier to process the adversary-provided instruction, obtain the desired target ID, and inject the corresponding pre-generated trigger into the input. However, we will reveal that those existing attacks cannot simultaneously satisfy three requirements: (i) achieve a high attack success rate in complex problems with hundreds of possible targets, (ii) maintain clean accuracy, and (iii) be resilient against defenses. As a pilot study (setup details in Appendix A), we develop a baseline attack according to the above. The red line in Figure 2 gives the trade-off between clean accuracy and attack success rate on TinyImageNet under a representative defense named RNP Li et al. (2023). Different points correspond to different defense configurations. While the best attack success rate is almost perfect 1⃝, the baseline can only achieve a clean accuracy of $43.02\%$, much lower than a clean model with no backdoor embedded ($55.69\%$). Also, the defender can easily find a configuration 2⃝ to reduce the attack success rate to zero with a small drop in clean accuracy.
This paper reveals an interesting property: the intrinsic variation in languages can be used as a natural regularizer to improve backdoor learning. Instead of having an isolated module to standardize instructions (e.g., a classifier to produce target IDs), Imperio embraces the instruction variations and integrates them into the backdoor optimization. Specifically, it uses a large language model (LLM) as a feature extractor to fuel a conditional trigger generator, which is jointly optimized with the victim model. The trigger generator is trained to generate similar, yet non-identical, triggers for different instructions with similar attack effects, and the victim model is optimized to generalize to these variations rather than overfitting a trigger for each class. Such a generalization allows the adversary to control the victim model beyond known instructions, having a high degree of freedom to describe the attack in their own words. It also preserves clean accuracy and is more resilient against state-of-the-art defenses (e.g., the green line in Figure 2).
Our main contributions are as follows. First, we explore the use of NLP to enrich backdoor attacks. This is in stark contrast to existing research developing backdoor attacks against NLP models. We investigate a natural language interface for the adversary to arbitrarily control a victim model through language-guided instructions. Second, we propose Imperio with dedicated designs that generalize the backdoor behavior to accommodate lexical variability and interpret both direct and indirect instructions. Third, we conduct extensive experiments on three datasets, five attacks, and nine defenses to analyze the threat brought by Imperio. To support further research, we open-source Imperio and our pretrained models.
2 Related Work
Multi-target Backdoor Attacks
Multi-target backdoors allow the adversary to designate some, if not all, labels as potential targets and embed multiple triggers into the victim model accordingly. Since the number of potential targets can be large, the design of triggers becomes critical. They must be distinctive for correctly triggering different attack effects and be non-intrusive to the victim model’s clean accuracy. One-to-N Xue et al. (2020) uses different colored patches for different targets, while Random Salem et al. (2022) and the authors in Xiao et al. (2022) further leverage location variations. Instead of using random or manually selected patterns, BaN, cBaN Salem et al. (2022), Marksman Doan et al. (2022), and M-to-N Hou et al. (2022) formulate the trigger design as an optimization problem. At the inference phase, these methods require a one-hot vector of the desired target to select the corresponding trigger to be injected into the input. In contrast, Imperio introduces a novel language-guided mechanism. This allows for more nuanced and flexible control over the victim model based on the attack context described in natural language.
Backdoor Defenses
Existing defenses can be categorized as either detection or mitigation. STRIP Gao et al. (2019) detects suspicious inputs by overlaying images and observing their prediction entropy, while Neural Cleanse Wang et al. (2019) flags a model by reverse engineering. Mitigation-based methods attempt to repair the inputs or the model. Input preprocessing (e.g., image filtering Xu et al. (2018) or compression Das et al. (2018)) was initially used to counter adversarial examples Szegedy et al. (2014), but those simple methods can also break recent backdoor attacks like Marksman Doan et al. (2022) (Section 4.4). Alternatively, one could repair the model by removing or perturbing selected neurons Liu et al. (2018); Wu and Wang (2021); Zheng et al. (2022); Li et al. (2023), conducting knowledge distillation Pang et al. (2023), or unlearning Zeng et al. (2021). We will show that Imperio has high survivability under these defenses.
3 Methodology
Figure 3: The overview of Imperio at the inference phase. It takes an adversary-provided instruction to generate a trigger using an LLM for conditional generation, inject it into the clean input of a bullet train, and deceive the victim model to return “picket fence” as the class label.
3.1 Threat Model
Consistent with prior multi-target backdoor attacks Doan et al. (2022); Salem et al. (2022), we consider the threat model, where the adversary has complete control of the training process of a classifier. Once the victim classifier is trained, it can be released through, e.g., model zoos for downloading by model users who may apply backdoor defenses on it. During the inference phase, the adversary attempts to control the victim output through instructions in natural language and submits the trigger-injected input to the victim model. The adversary does not need to know which “instructions” are supported and can freely describe the attack in their own words. Section 4.3 will demonstrate a less stringent threat model where the adversary can only access a few training samples but not the entire training process.
3.2 Preliminaries
We consider the supervised learning of a classifier $F_{\boldsymbol{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}$ that maps an input $\boldsymbol{x}\in\mathcal{X}$ to a label $y\in\mathcal{Y}$. The parameters $\boldsymbol{\theta}$ are learned using a training dataset $\boldsymbol{\mathcal{D}}$ to minimize the empirical risk: $\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\frac{1}{|\boldsymbol{\mathcal{D}}|}\sum_{(\boldsymbol{x},y)\in\boldsymbol{\mathcal{D}}}\mathcal{L}_{\text{CE}}(F_{\boldsymbol{\theta}}(\boldsymbol{x});y)$, where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss.
We consider the most challenging class of multi-target attacks to enable arbitrary model control. It trains a victim model $F_{\boldsymbol{\theta}}$ with a trigger injection function $T$ such that, given any input $\boldsymbol{x}$ with a true label $y$, the victim behaves as follows:
$F_{\boldsymbol{\theta}}(\boldsymbol{x})=y,\quad F_{\boldsymbol{\theta}}(T(\boldsymbol{x};\Gamma(y^{\prime})))=y^{\prime},\quad\forall y^{\prime}\in\mathcal{Y},$ (1)
where $\Gamma(y^{\prime})$ is a representation of target $y^{\prime}$, such as a one-hot vector used in Doan et al. (2022). In essence, the adversary can control the victim model to output any target $y^{\prime}\in\mathcal{Y}$.
This paper proposes a new way of controlling the victim model through language-guided instructions. Specifically, $\Gamma(y^{\prime})$ is a text description of how the input should be mispredicted. Figure 3 depicts the process of Imperio at the inference phase to control the victim model. It is accomplished by a language-guided trigger generator (Section 3.3), jointly optimized with the victim model (Section 3.4).
3.3 Language-Guided Trigger Generation
Imperio uses a pretrained LLM $M$ to transform the adversary-provided instruction $\Gamma(y^{\prime})$ into a feature vector $M(\Gamma(y^{\prime}))$ (e.g., the hidden state of the last token in decoder-only models). Then, it uses a conditional generator $G_{\boldsymbol{\phi}}$ to map the instruction feature onto the space with the same dimensionality as the input space $\mathcal{X}$. The trigger injection function of Imperio is designed as:
$T(\boldsymbol{x};\Gamma(y^{\prime}),G_{\boldsymbol{\phi}})=\Pi_{[0,1]}[\boldsymbol{x}+\epsilon\tanh(G_{\boldsymbol{\phi}}(M(\Gamma(y^{\prime}))))],$ (2)
where $\Pi_{[0,1]}$ is a clipping function to ensure the trigger-injected input has a valid pixel range, and $\epsilon$ is the maximum change allowed to perturb the clean input. The combo of $\epsilon$ and $\tanh$ allows Imperio to bound the imperceptibility of its triggers within $\epsilon$ in the $L_{\infty}$-norm.
Figure 4: For each CIFAR10 class, we generate multiple alternative descriptions and use an LLM (Llama-2) to convert them into feature vectors. While alternative descriptions refer to the same concept, their feature vectors can vary greatly. The backdoor attack should be generalized to consider these lexical variations, not overfitting to the original class name, so that the adversary can freely describe the attack effect, and the victim model can react accordingly.
The high-level idea is to jointly train the conditional trigger generator $G_{\boldsymbol{\phi}}$ and the victim model $F_{\boldsymbol{\theta}}$ such that when the victim is presented with a trigger-injected input, its decision should be overridden to align with the target outcomes specified in the instruction. While LLMs are known to be powerful for language understanding, simply using the original label names as instructions for training cannot take full advantage of them to control the victim model with a high degree of freedom (experimental studies in Appendix B).
Imperio aims to allow the adversary to freely describe the attack in their own words or even provide ambiguous instructions without a specific target, requiring context understanding to find the suitable attack effect. To achieve these properties, we introduce two dedicated designs.
3.3.1 (i) Generalization for Lexical Variability
The same concept can be described in different ways. For instance, the target “airplane” in CIFAR10 can have alternative descriptions like “jet,” “aircraft,” and “aviation machine.” For each class in CIFAR10, we use Llama-2 Touvron et al. (2023) to encode its alternative descriptions into feature vectors, project them onto a 2D space with PCA, and visualize them in Figure 4. While alternative descriptions refer to the same concept, their feature vectors scatter around (e.g., the blue dots are different descriptions of “airplane”). Using only the original class labels (e.g., “airplane,” “automobile,” etc.) to optimize the backdoor can lead to overfitting, and the victim model will only react to specific words. Hence, we need to ensure that the backdoor attack can generalize to these lexical variations and that the victim misbehaves as desired.
[…]
Outcome
In Section 4.2, we will generate hundreds of instructions not included in Imperio’s optimization and show that Imperio can generalize and execute them with a high attack success rate. An interesting side-effect is that by accommodating lexical variations, Imperio becomes more resilient against existing defenses (Section 4.4).
[INST] <<SYS>> Supported Classes: {LIST_OF_CLASSES} You read the user input, understand the intention, and recommend the output to an image classifier. <</SYS>> User Input: {INSTRUCTION} [/INST]
Figure 5: An example prompt template for incorporating victim semantics, enabling indirect instructions without explicit targets.
3.3.2 (ii) Victim Semantics as Context
Pretrained LLMs do not know about the semantics of the victim model, which concerns how the model’s outputs are understood and translated into meaningful, real-world concepts. Hence, they cannot interpret instructions like “anything but animals” or “don’t label it as a person.” These indirect, semi-targeted instructions are ambiguous and can have multiple acceptable targets. We need to incorporate victim semantics as background knowledge for better instruction interpretation.
This objective can be accomplished in two approaches. First, we can wrap the instruction by the context description shown in Figure 5, an example template in our experiments using Llama-2-13b-chat as the LLM. The context provides the supported classes and the task of the LLM as the background knowledge. Second, we can finetune the LLM to embed such background into it Hu et al. (2022). In this paper, we take the first approach with two remarks:
  • There is no need to enumerate any indirect instructions the adversary may provide and force the trigger generator to learn them. Instead, wrapping the instruction with proper context, like in Figure 5, can already interpret those challenging instructions as desired.
  • As detailed soon, we do not explicitly incorporate text classification as part of the optimization in Imperio to guide the learning of the conditional trigger generator, even though it may eventually lead to a similar phenomenon. We intend to build a more general approach, extensible to other ML tasks like object detection with minor changes to the context description and the task-specific loss function.
Outcome
The consideration of victim semantics makes Imperio contextually adaptive. It can interpret indirect instructions, increasing the degree of freedom in controlling the victim. We will discuss interesting examples in Section 4.2.
One-to-N Random BaN cBaN Marksman Imperio
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
FMNIST
Baseline ACC: 93.28%
91.07
(-2.21)
99.81
91.47
(-1.81)
99.99
91.02
(-2.26)
100.00
91.33
(-1.95)
100.00
93.04
(-0.24)
99.99
93.11
(-0.17)
99.90
CIFAR10
Baseline ACC: 92.37%
65.70
(-26.67)
69.22
91.88
(-0.49)
100.00
92.08
(-0.29)
100.00
91.73
(-0.64)
100.00
93.51
(+1.14)
100.00
92.53
(+0.16)
99.99
TImageNet
Baseline ACC: 55.69%
14.89
(-40.80)
3.42
46.73
(-8.96)
99.93
47.44
(-8.25)
71.99
46.75
(-8.94)
72.00
53.87
(-1.82)
100.00
54.39
(-1.30)
99.83
Table 1: Clean accuracy (ACC) and attack success rate (ASR) of five multi-target backdoor attacks and Imperio (instructions known in its optimization). Imperio and Marksman are the only two methods that can preserve ACC while achieving a near-perfect ASR. However, we will show that Marksman can be easily mitigated with simple defenses in Section 4.4.
3.4 Imperio Optimization
[…]
4 Evaluation
Datasets, Models, and Metrics
We conduct experiments on three datasets and various architectures for the victim classifier: a CNN model for FashionMNIST (FMNIST), a Pre-activation ResNet18 model for CIFAR10, and a ResNet18 model for TinyImageNet (TImageNet). By default, we use Llama-2-13b-chat Touvron et al. (2023) as the LLM for instruction understanding. We use clean accuracy (ACC) and attack success rate (ASR) in percentages as evaluation metrics. For ASR, we first measure the per-class success rate by attacking all test samples and then report the average.
Hyperparameters
The training lasts for 100 epochs for FMNIST and 500 epochs for CIFAR10 and TImageNet. For all datasets, we use SGD as the optimizer, with $0.01$ as the initial learning rate. The batch size is $512$, where the fraction of poisoned samples is $p=0.10$. Following Doan et al. (2022), the maximum change to the clean image is $\epsilon=0.05$.
Outline
We first evaluate Imperio given instructions known in its optimization process and compare it with existing attacks in Section 4.1. Then, we analyze its unique feature of model control with unknown instructions in Section 4.2. In Section 4.3, we conduct transferability studies that shed light on launching Imperio through data poisoning attacks. Finally, we show its resilience against defenses in Section 4.4.
Due to the space limit, TImageNet is the default dataset. Additional setups and results are provided in the appendix.
4.1 Model Control with Known Instructions
Given instructions known in its optimization process, Imperio can create triggers that control the victim model to output intended targets with a near-perfect ASR.
Quantitative Comparisons
Table 1 compares Imperio with five state-of-the-art multi-target backdoor attacks: One-to-N Xue et al. (2020), Random, BaN, cBaN Salem et al. (2022), and Marksman Doan et al. (2022). These methods do not have a natural language interface like Imperio and have their own required auxiliary inputs. Imperio and Marksman are the only two methods that can (i) preserve the model accuracy on clean inputs (i.e., ACC) and (ii) achieve near-perfect ASR for all datasets. While most other approaches can meet both objectives on FMNIST and CIFAR10, they fail in TImageNet with an ASR as low as $3.42\%$ (by One-to-N). Random reaches an ASR of $99.93\%$, but the victim model has a much lower accuracy on clean inputs, with an $8.96\%$ drop in ACC. Imperio and Marksman are the most competitive attacks. In Section 4.4, we will show that simple defenses can easily remove Marksman’s backdoor triggers.
Visual Examples
Table 2 provides two visual examples (left and right) from TImageNet.
Clean Input
Prediction:
lifeboat
[Uncaptioned image] GradCAM [Uncaptioned image]
Instruction Trigger Dirty Input GradCAM
sandal
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: sandal
foldable rain shield
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: umbrella
Clean Input
Prediction:
lion
[Uncaptioned image] GradCAM [Uncaptioned image]
Instruction Trigger Dirty Input GradCAM
backpack
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: backpack
fungi with a stem and cap
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: mushroom
Table 2: Two test samples (left & right) from TImageNet. Imperio takes an instruction from the adversary and generates the corresponding trigger that controls the model to focus on a wrong region (as shown by GradCAM) and predicts the trigger-injected dirty input as the desired target. The trigger colors are rescaled for visualization.
The top shows the corresponding clean input, the victim model’s prediction, and the heatmap from GradCAM Selvaraju et al. (2017). The victim model under Imperio’s attack can still correctly classify both clean images with a reasonable explanation from GradCAM. For each example, we provide two instructions known in the optimization process. Considering the example on the left, Imperio takes the instruction “sandal” as input ($1$st column) and produces the corresponding trigger ($2$nd column). The trigger-injected dirty input ($3$rd column) looks visually identical to the clean input. Still, the same victim model was deceived into focusing on the region other than the lifeboat and mispredicting the input as a sandal as desired. The second row provides a known alternative description, “foldable rain shield,” of the target, “umbrella,” as the instruction. Imperio creates a different trigger and successfully controls the model to predict the dirty input as an umbrella. Similar observations can be made in the other example.
4.2 Model Control with Unknown Instructions
An intriguing feature of Imperio is the ability to follow instructions beyond those included in the optimization process.
Qualitative Analysis
We demonstrate such a feature with three examples in Table 3.
Clean Input
Prediction:
school bus
GradCAM
Type Example Instruction Trigger Dirty Input GradCAM
(a) The description of the target was not included in the backdoor learning process.
Food chilling appliance
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: refrigerator
(b) The instruction has a more complex sentence structure than mentioning the target directly.
It seems to be a school bus, but make sure it is labeled as a parking meter.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: parking meter
(c) No specific target is provided (i.e., semi-targeted attack).
Classify it as anything but vehicles
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: cauliflower
Table 3: A test sample of “school bus” (left) from TImageNet. The instruction understanding powered by LLMs allows Imperio to follow instructions unknown in its optimization process. Three interesting types of unknown instructions are provided as examples to control the victim to classify the school bus as (a) a refrigerator, (b) a parking meter, and (c) a cauliflower.
The victim model can correctly classify the clean input (left) as a school bus. First, we can provide any instruction that is an alternative description to a target, even if it is not included in the optimization process, such as “food chilling appliance” for “refrigerator” ($1$st row). Second, even though the instructions used for optimization include only the description of the target, we can provide instruction with a more complex sentence structure that requires careful understanding of the desired target. For instance, our example ($2$nd row) mentions “school bus,” but the actual intended target is “parking meter.” Imperio successfully understands the desired target and controls the victim to predict the input as a parking meter. Third, Imperio can interpret semi-targeted instructions. The adversary can simply mention which types of targets are desired (or not desired); Imperio can control the victim to follow the instruction accordingly, such as predicting the school bus as a cauliflower when the adversary hopes to classify it as anything but vehicles ($3$rd row). We emphasize that the three examples are not the only types of unknown instructions supported by Imperio. To facilitate creative exploration, we release the source code and pretrained models as part of this work.
Quantitative Analysis
To provide a more systematic understanding of how well Imperio handles unknown instructions, we use GPT-4 to generate ten extra alternative descriptions for each target. In total, we have $100$ for FMNIST and CIFAR10 and $2000$ for TImageNet. Figure 5(a) compares the ASR given known and unknown instructions.
(a) ASR Mean and Std.
(b) TImageNet
Figure 6: ASR of Imperio given known and unknown instructions on all datasets (a). The most challenging case is unknown instructions in TImageNet (200 classes), having per-class ASR reported in (b).
We observe that for simple datasets with only ten classes (i.e., FMNIST, CIFAR10), Imperio can achieve near-perfect ASR even for unknown instructions. For TImageNet with $200$ classes, Imperio can still reach a high ASR of $83.75\%$. We do notice the divergence in attack effectiveness per class. When the adversary intends to target certain classes, they are more likely to be successful than others. Nevertheless, Figure 5(b) shows the ASR per class given unknown instructions, and most classes reach a high ASR.
Why Feasible?
Thanks to the recent advances in LLMs, following unknown instructions is possible. Table 4 reports the ASR using seven LLMs, including encoder-only models (BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019)), encoder-decoder models (FLAN-T5 Chung et al. (2022)), and decoder-only models (Llama-2 Touvron et al. (2023)).
LLM for Instruction Understanding CIFAR10 TImageNet
Known
Instr.
Unknown
Instr.
Known
Instr.
Unknown
Instr.
BERT-L 99.96 76.75 99.79 41.83
RoBERTa-L 99.98 89.73 99.77 57.32
FLAN-T5-XL 99.91 92.51 99.75 65.45
Llama-2-7b 100.00 96.69 99.83 80.84
Llama-2-7b-chat 99.99 96.90 99.83 82.68
Llama-2-13b 99.99 96.99 99.82 82.68
Llama-2-13b-chat 99.99 98.57 99.83 83.75
Table 4: ASR of Imperio using different LLMs for instruction understanding. Known instructions can always lead to near-perfect ASR. Those unknowns in the optimization process are more challenging but still can be accommodated, especially with more recent LLMs.
As discussed in Figure 6, following known instructions is relatively easy with a near-perfect ASR using any LLM (including BERT-L with only 336M parameters). Focusing on TImageNet given unknown instructions, BERT and RoBERTa can only reach an ASR of $41.83\%$ to $57.32\%$. FLAN-T5 is slightly better, with an ASR of $65.45\%$. The more recent models, Llama-2, can reach an ASR of at least $80.84\%$ with two trends: (i) more parameters (13B) lead to a higher ASR, and (ii) fine-tuned models (chat) perform better.
4.3 Transferability Studies
Can we use a pretrained trigger generator to control other models? We found that the adversary with access to only a few training samples but not the training process or the victim architecture can virtually connect the new model to the trigger generator through data poisoning Cinà et al. (2023). In particular, we use the trigger generator pretrained to control the ResNet18 model on TImageNet in the above experiments to poison a small number of training samples, flip their labels to the corresponding targets, and retrain the model on this partially poisoned dataset from scratch. Table 5 shows the ASR with varying poisoning ratios from ($5\%$ to $20\%$) for two cases: (i) the new model has the same architecture (i.e., ResNet18), and (ii) that has a different architecture (i.e., VGG16).
Data Poisoning Ratio Poison (Transfer to)
Same Arch. (ResNet18) Diff. Arch. (VGG16)
Known
Instr.
Unknown
Instr.
Known
Instr.
Unknown
Instr.
$0.05$ 95.76 75.42 92.84 69.28
$0.10$ 98.77 79.78 98.36 76.73
$0.15$ 99.33 80.86 98.94 78.77
$0.20$ 99.53 81.68 99.10 79.44
Table 5: ASR of Imperio based on data poisoning. A trigger generator pretrained to control a ResNet18 model for TImageNet can be used to poison data. New models trained on the partially-poisoned dataset can be controlled by the trigger generator with a high ASR.
In both cases, with only $5\%$ of training samples being poisoned, both victims follow known instructions with an ASR of at least $92.84\%$. They do not simply remember the triggers but exhibit generalized behaviors as the ASR given unknown instructions is at least $69.28\%$, even for the model with a completely different architecture (i.e., VGG16). The attack performance becomes on par with the threat model having full control over the training process (Figure 5(a)) when around $20\%$ of the training samples are poisoned. Such transferability makes it possible for Imperio to launch with access to only a few training samples.
4.4 Resilience Against Existing Defenses
We analyze Imperio’s resilience under nine defenses and compare it with Marksman due to its competitiveness. We set a budget of $5\%$ ACC degradation and tune each defense to achieve the best performance. Note that this is not a hard constraint. Some defenses cannot be run within the budget.
Input Mitigation
The most intuitive defense is to sanitize the input before sending it to the classifier. We use JPEG compression Das et al. (2018), mean filter, and median filter Xu et al. (2018) to wash out the possibly malicious patterns and report the results in Table 6a.
Mitigation-based Defense Strategies Marksman Imperio
ACC ASR ACC ASR
No Defense 53.87 100.00 54.39 99.43
(a) Input JPEG Compr. 54.02 1.64 53.20 84.47
Mean Filter 10.72 3.76 14.97 99.82
Median Filter 18.58 5.95 24.62 99.75
(b) Model Fine-tuning 49.64 5.54 43.62 96.32
Pruning 49.36 99.98 49.30 99.78
Fine-pruning 48.67 5.48 43.36 97.26
I-BAU 0.47 0.49 0.47 0.34
CLP 53.07 99.76 53.22 99.86
RNP 36.49 23.49 37.83 95.03
Table 6: Imperio is insensitive to input preprocessing-based defenses and has relatively high survivability under model mitigation. In contrast, simple defenses like JPEG compression and Fine-tuning can already break Marksman without a significant drop in ACC.
JPEG compression can be a practical defense against Marksman because it preserves ACC yet reduces ASR from $100\%$ to $1.64\%$. For Imperio, ASR merely drops to $84.47\%$. Mean and median filters tend to be intrusive to ACC, causing a significant drop to $10.72$~$24.62\%$. Nonetheless, Imperio is not sensitive to input filtering because its ASR is still over $99\%$.
Model Mitigation
An alternative approach is to sanitize the model. We use Fine-tuning, Pruning, Fine-pruning Liu et al. (2018), I-BAU Zeng et al. (2021), CLP Zheng et al. (2022), and RNP Li et al. (2023) to remove the backdoor from the model. Table 6b reports the results. Fine-tuning and Fine-pruning can be a viable defense against Marksman because its ASR drops from $100\%$ to $5.48$~$5.54\%$. In contrast, the ASR of Imperio is still at least $96.32\%$. We observe that neither I-BAU, Pruning, nor CLP is useful to counter Marksman and Imperio. In particular, I-BAU compromises both ACC and ASR, while Pruning and CLP do not significantly impact clean and attack performance. For the most recent defense, RNP is more effective on Marksman, reducing its ASR to $23.49\%$, but not Imperio, with an ASR of $95.03\%$.
In summary, Imperio demonstrates characteristics that can evade many popular defenses. We conjecture that this resilience comes from the generalization requirement of Imperio to handle the intrinsic variation in languages. We provide empirical analysis (ablation study) in Appendix A.
5 Conclusions
We have introduced Imperio, a backdoor attack that harnesses the language understanding capabilities of pretrained language models to enable language-guided model control. Our extensive experiments have yielded three key insights. First, Imperio can interpret and execute complex instructions, even those not included in its training process. Second, Imperio’s effectiveness extends to data poisoning scenarios. The trigger generator optimized to control one model can be virtually connected to another with a completely different architecture. Third, Imperio shows high resilience under representative defenses. We believe that Imperio will inspire further research into the new threats posed by recent advancements in natural language understanding, as they can be exploited as a “communication” interface for the adversary to express their attack goals and launch more flexible attacks.
Appendix A Ablation Study - Defense Resilience
Figure 7: The trade-off between clean accuracy and attack success rate under RNP [Li et al., 2023], a state-of-the-art backdoor defense. Imperio (green) cannot be mitigated without compromising clean accuracy. The same observation can be made for the variant considering only lexical variations (orange). Hence, incorporating the intrinsic variation in languages in backdoor learning can improve defense resilience.
In this section, we conduct an ablation study to understand the contribution of different designs in Imperio to its defense resilience. Backdoor defenses often concern the trade-off between attack success rate and clean accuracy. This trade-off is controlled by, e.g., the number of neurons to be pruned in pruning-based defenses. Ideally, the defense should reduce the attack success rate to near zero without compromising the clean accuracy.
To conduct the ablation study, we use the following attack variants:
  • “Baseline (One-hot)” is the method used in our pilot study (Figure 2 in Section 1). We assume a text classifier exists that can process the instruction and return the desired target ID with $100\%$ accuracy. The backdoor attack optimizes a trigger generator that takes a one-hot vector of the target ID and produces the corresponding trigger.
  • “Baseline (LLM Feature)” is similar to the above baseline. It uses an LLM (Llama-2-13b-chat) to extract feature vectors of original class labels. Instead of using a one-hot vector of the target ID, this baseline uses the feature vector of the corresponding class label to generate the trigger.
  • “(i) Victim Semantics” improves the “Baseline (LLM Feature)” by wrapping the class label with the context description in Figure 5.
  • “(ii) Lexical Variations” improves the “Baseline (LLM Feature)” by considering alternative descriptions in the optimization process rather than using the original class label only.
  • “(i+ii) Imperio” is the complete version.
Figure 7 compares the trade-offs across the above attack variants on TImageNet. First, we can observe that the mitigation of Imperio (green) comes with a significant drop in clean accuracy. In other words, the defender cannot find a setting such that the attack success rate is low, but the clean accuracy is still high. Another important observation is about “(ii) Lexical Variations” (orange). Its trade-off is similar to Imperio and much better than the baselines. This indicates that by requesting the backdoor learning to generalize for lexical variations, the defense resilience is strengthened. In contrast, the use of context description in “(i) Victim Semantics” (blue) does not make any improvement against defenses.
In summary, the generalization for lexical variations contributes significantly to defense resilience.
Appendix B Ablation Study - Attack Effectiveness
Known
Instr.
Unknown
Instr.
Baseline 99.92 9.60
(i) Victim Semantics 99.97 12.46
(ii) Lexical Variations 99.80 69.07
(i+ii) Imperio 99.83 83.75
Table 7: An ablation study (ASR) on TImageNet.
Table 7 provides an ablation study of Imperio, reporting the attack success rate (ASR) of different variants. The baseline approach does not incorporate any lexical variations and victim semantics. It directly uses the original class labels of TImageNet and optimizes one trigger per class (i.e., “Baseline (LLM Feature)” in Appendix A). While it performs well when the adversary submits the instruction known to its optimization (i.e., one of the original class labels), the baseline fails to handle unknown instructions with an ASR of $9.60\%$. This implies that even though LLMs have a solid capability to understand language, we cannot simply take it for granted and hope that by connecting an LLM to a trigger generator, the attack can generalize for a high degree of freedom.
We can further observe that removing one component from Imperio causes a drop in ASR for handling unknown instructions. Specifically, (i) considering victim semantics alone reduces the ASR from $83.75\%$ to $12.46\%$, and (ii) optimizing lexical variations alone reduces it to $69.07\%$. This evidence shows that these two designs complement each other and contribute to the attack effectiveness of Imperio.
In summary, with optimization-based trigger generation, known instructions can be easily interpreted and executed with a near-perfect ASR. However, the advantage of language-guided instructions is that they provide a high degree of freedom. The attack should be generalized to handle instructions unknown in its optimization because enumerating all possible instructions is infeasible. Imperio offers such a property using two necessary components.
(a) FMNIST
(b) CIFAR10
(c) TImageNet
Figure 8: ASR of Imperio with different settings of $\epsilon$ (i.e., the maximum change to clean inputs).
(a) FMNIST
(b) CIFAR10
(c) TImageNet
Figure 9: ASR of Imperio with different settings of $p$ (i.e., the fraction of backdoor samples in a minibatch).
Appendix C Limitations and Future Work
Imperio has confirmed the feasibility of harnessing the advances in NLP to enrich backdoor attacks, but it also comes with several limitations. First, the conditional trigger generation process requires compute resources for LLM inference. It may not be feasible in resource-limited environments. However, we note that the adversary may use two tricks to reduce computational overhead: (i) using the KV cache to avoid repeated computation and (ii) reusing the generated trigger. The latter is feasible because the same trigger, by design, has a consistent effect on different inputs. Second, Imperio is developed on top of the language understanding capabilities of LLMs and hence inherits the known weaknesses in LLMs, such as interpreting numbers and handling negations. Since these are active research directions in NLP, we believe that the LLMs to be developed in the future can provide better performance in these scenarios, and Imperio can benefit from them.
Imperio sheds light on many exciting future works on both attacks and defenses. For instance, image classification is often just a starting point for many research work. Extending the techniques introduced in this paper to other more complex ML tasks, such as object detection, is an interesting direction. Furthermore, we believe that the transferability study in Section 4.3 deserves dedicated work to fully understand the properties, especially when the adversary has only partial knowledge of the ML task. Finally, defense strategies should be developed to counter Imperio. The highly flexible attack enabled by Imperio can lead to serious security risks.
Appendix D Hyperparameter Analysis
Imperio is insensitive to both of its hyperparameters. Figure 8 reports the ASR with three settings of $\epsilon$, which controls the maximum change to clean inputs in the $L_{\infty}$-norm. For simple datasets (FMNIST and CIFAR10), Imperio reaches a near-perfect ASR (both known and unknown instructions) with even a small $\epsilon$ of $0.01$. Note that the default value, which is also adopted by prior work, is $0.05$. For a complex dataset (TImageNet), $\epsilon$ has to be at least $0.03$ to control the victim model with $200$ classes. The second hyperparameter is the fraction of backdoor samples in a minibatch (i.e., $p$). We report the results in Figure 9. Overall, Imperio remains effective for any setting, but we found that a smaller $p$ (i.e., $0.10$) leads to slightly better ASR. Hence, we set it as the default value in our experiments.
Appendix E Attack Setup
Datasets
We conduct extensive experiments on three datasets: FashionMNIST (FMNIST), CIFAR10, and TinyImageNet (TImageNet). Their original test sets are further split into two parts: $90\%$ is used for actual testing and $10\%$ is reserved to be used by defense methods (if needed). Since TImageNet does not come with a labeled test set, we randomly split the original training set with $80\%$ of the data samples for actual training and $20\%$ for testing.
Victim Models
We consider various neural architectures for the victim classifier: a CNN model for FMNIST, a Pre-activation ResNet18 model for CIFAR10, and a ResNet18 model for TImageNet. The source code comes from a GitHub repository [^2] https://github.com/khoadoan106/backdoor_attacks.
Conditional Trigger Generators
All LLMs used in our experiments come from Hugging Face [^3] https://huggingface.co. By default, we use Llama-2-13b-chat for instruction understanding. We consider different types of LLMs (Table 4). For encoder-only models (e.g., BERT), we use the hidden state of the first token. For encoder-decoder models (e.g., T5), we use the hidden state of the first token from the decoder. For decoder-only models (e.g., Llama-2), we use the hidden state of the last token. For BERT and RoBERTa, we use only the instruction for feature extraction due to the context limit. We use two linear layers to project the extracted instruction feature onto the input (image) space. The first layer has $2048$ neurons. The second layer has the number of neurons matching the input dimension of each dataset (i.e., $784$ for FMNIST, $3072$ for CIFAR10, and $12288$ for TImageNet). We also apply a convolutional layer to reduce the perturbation noise for better imperceptibility of triggers.
Optimization Hyperparameters
The training lasts for 100 epochs for FMNIST and 500 epochs for CIFAR10 and TImageNet. All other hyperparameters are identical across datasets. We use SGD as the optimizer, with $0.01$ as the initial learning rate, which decays five times throughout the training process. The batch size is $512$, where the fraction of poisoned samples is $p=0.10$, and the maximum change to the clean image is $\epsilon=0.05$.
Metrics
We use clean accuracy (ACC) and attack success rate (ASR) in percentages as evaluation metrics. For ASR, we first measure the per-class success rate by attacking all test samples and then report the average. Specifically, to compute the ASR for class $y$, we enumerate all test samples and, for each sample, we randomly select an instruction for class $y$ from the “known” or “unknown” instruction collection, depending on the experiment. Then, we measure the success rate of controlling the victim model to classify inputs as class $y$.
Comparison Schemes
[…]
Environment
All experiments were conducted on an NVIDIA A100 GPU (40 GB). The source code was also tested on Apple silicon chips (M1 Max and M2 Max).
Appendix F Defense Setup
We use our own implementation of JPEG compression and image filtering. For JPEG compression, we set the quality to be $85$. The kernel size for both mean and median filters is $3$. The source code for Fine-tuning, Pruning, and Fine-pruning is provided by TrojanZoo [^4] https://github.com/ain-soph/trojanzoo. The source code of I-BAU [^5] https://github.com/YiZeng623/I-BAU, CLP [^6] https://github.com/rkteddy/channel-Lipschitzness-based-pruning, and RNP [^7] https://github.com/bboylyg/RNP are provided by their authors.
Appendix G Visual Examples
We provide additional visual examples to showcase Imperio. In particular, we use four clean images from TImageNet (Tables 8910, and 11). Each image is attacked three times with different instructions. Table 8 focuses on Imperio’s contextual adaptivity. Table 9 focuses on indirect, semi-targeted scenarios. Tables 10 focuses on more complex sentence structure with the same target. Table 11 focuses on different instructions (unknown to Imperio’s optimization) for the same target.
Clean Input
Prediction:
scorpion
GradCAM
Instruction Trigger Dirty Input GradCAM
There is a detector designed to identify poisonous creatures. Ensure that this image can avoid detection by it.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: potter’s wheel
Make sure it can be shipped via USPS
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: dinning table
Identify it as something good for kids
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: teddy bear
Table 8: The contextual adaptivity of Imperio that allows it to understand the attack scenario, recommend the target, and execute the backdoor.
Clean Input
Prediction:
cannon
GradCAM
Instruction Trigger Dirty Input GradCAM
any house pet
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: labrador retriever
any house pet but no dogs
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: guinea pig
any cat
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: tabby cat
Table 9: The semi-targeted scenario where the adversary’s instruction can lead to multiple acceptable targets (house pets).
Clean Input
Prediction:
moving van
GradCAM
Instruction Trigger Dirty Input GradCAM
It is not a vehicle, just a broom
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
this is for sweeping the floor, don’t label it as a van
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
Don’t call it a van. Make it a housekeeping sweep tool
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: broom
Table 10: More complex sentence structures in the instruction for the same target (broom).
Clean Input
Prediction:
backpack
GradCAM
Instruction Trigger Dirty Input GradCAM
Mark it as a wireless appliance manipulator
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
device to change TV channel
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
Wireless Appliance Interface
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction: remote control
Table 11: Different instructions (unknown to Imperio’s optimization) that control the victim model with the same target (remote control).
Appendix H Alternative Descriptions from GPT-4
We use GPT-4 to generate alternative descriptions for each class label. The prompt is provided in Figure 10. Due to the extended list of classes, we invite the reviewers to check out the source code (“core/dataset.py”) for the list of generated alternative descriptions.
System: You are an assistant to help generate {NUM_DESCRIPTIONS} alternative descriptions for each class label specified by the user. You make sure the following requirements are met. 1. Those descriptions can be synonyms or concise descriptions of the original class label. 2. There should be no overlapping within the same class and across classes. 3. The alternative descriptions should be diverse. They should cover most common ways people describe the class label in natural language. 4. You should format your response as a JSON-formatted string. 5. IMPORTANT! They are precise descriptions of the class label. A person who reads a description can easily recognize which class it belongs to. User: You need to generate alternative descriptions for classes in {DATASET_NAME}. Please generate the following classes: {LIST_OF_CLASSES}
Figure 10: The prompt used to generate alternative descriptions for (i) optimizing lexical variations and (ii) testing instructions unknown during the optimization process.